in this formula c= $5.25+ $2.50h c represents the toal charge for babysitting and h represents the number of hours the child is kept. Which is the amoubt joe should pay when his child is at the babysitting service for 4hrs?
Answer by Theo(13342)

joe pays 5.25 + 4 * 2.50 = 5.25 + 10.00 = 15.25

h became 4.